OpenStudy (anonymous):

If m men can paint a house in d days, how many days will it take m+2 men to paint the same house? A. d+2 B. d-2 C. (m+2)/md D. md/(m+2) E. (md+2d)/m

OpenStudy (anonymous):

\[\frac{ m }{ d } = \frac{ m+2 }{ x }\]\[mx = (m+2)d\]\[x = \frac{ md + 2d }{ m }\]

OpenStudy (anonymous):

E.